# On the Algorithmic Bias of Aligning Large Language Models with RLHF: Preference Collapse and Matching Regularization

### v1

@@ 5.2 Extreme Case: Preference Collapse in KL RLHF
Preference collapse as an extreme form of the algorithmic bias, which is discussed in Section 5.1, arises due to the use of the reference model and continues to occur even if the KL divergence is replaced by any $f$-divergence. To see this, note that (5.2) remains true if $p_{\textnormal{reward}}$ is replaced by $p^{f}_{\textnormal{reward}}$. This type of bias inherits from the reference model which is, by nature, a biased LLM since it is not aligned yet.
6 Experiments
In this section, we demonstrate the effectiveness of conditional PM RLHF through a comparison with KL RLHF. The experiments were conducted on a computing infrastructure equipped with four A100 GPUs, each with 80GB of memory. Throughout the experimental process, each stage of RLHF follows the standardized configurations established in the DeepSpeed-Chat framework (Yao et al., 2023). [^11] The code is available at https://github.com/JiancongXiao/PM_RLHF.
Reward Model.
The reward model is trained on a dataset of human preferences. Specifically, we utilize the full-hh-rlhf dataset [^12] Dataset is availble at https://huggingface.co/datasets/Dahoas/full-hh-rlhf., available from Hugging Face, to fine-tune our models. This dataset comprises 112,000 training examples and 12,500 test cases. In the second stage, the reward model is trained using a network architecture identical to Llama-2-7B or OPT-1.3B, but with a different prediction head that outputs a scalar reward value. The training process involves minimizing the loss function defined in (2.2), with early stopping applied to prevent over-optimization (Schulman, 2023). It is worth noting that our theory requires $\beta=1$, and our experiments focus on this case. However, notice that (2.3) is equivalent to
6.1 Evaluation Metrics
Preference Matching Divergence.
Here the models can be the KL RLHF–aligned or conditional PM RLHF–aligned models. Then, the instance preference matching divergence given $x$ is defined by the KL divergence between the above two distributions, i.e.,
$\text{Instance PM Divergence}=D_{\textnormal{KL}}(p_{\text{llm}}(y|y_{1},y_{2},x)\|p_{\textnormal{reward}}(y|y_{1},y_{2},x)).$
The aggregate preference matching divergence is determined by calculating the expected value of the instance preference divergence over all instances of $x$ drawn from the test dataset.
PM $\text{PM Divergence}=\mathbb{E}_{x}\text{Instance PM Divergence}(x).$
A lower PM divergence suggests better performance in preference matching. When the PM divergence approaches 0, the preference of the LLMs is perfectly aligned with the preference induced by the reward models.
  • Perplexity. The perplexity (Jelinek et al., 1977) of an LLM for generating a response $y$ given an input $x$ can be characterized by the output probability $\pi_{\phi}(y|x)$. It is important to note that $\pi_{\phi}(y|x)$ represents the product of the probabilities of each token in $y$, making $\pi_{\phi}(y|x)$ not directly comparable across responses of different lengths. To address this, one can take the geometric mean of $\pi_{\phi}(y|x)$ for each token and then consider its inverse. This gives us the following standard definition of perplexity for a language model over a dataset $D$. Let $\pi_{\phi}(y|x)=p_{1}\times p_{2}\times\ldots\times p_{\text{Length}}(y)$, where $p_{i}$ is the probability of token $T_{i}$ given all the previous tokens, we have
    $\text{Perplexity}=\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}\big{[}p_{1}\cdot p_{2}\ldots p_{\text{Length}}(y)\big{]}^{-1/\text{Length}(y)}.$
  • Entropy. The entropy measures the diversity of the outputs. It is defined as
    $\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}-\log\pi_{\phi}(y|x).$
  • KL divergence. This metric measures the KL divergence between the trained model and the reference model. It is defined as
    $\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}\log\frac{\pi_{\phi}(y|x)}{\pi_{\textnormal{ref}}(y|x)}.$
6.2 Fine-Tuning LLMs
We will begin by presenting our experiments on the Llama-2-7B model, followed by the results obtained from the OPT-1.3B model.
Experiments on Preference Matching Divergence.
We begin by examining the preference matching divergence. Figure 5 presents the PM divergences in panels (a), (b), and (c) for $\beta$ values of 0.1, 0.5, and 1, respectively. Table 2 provides detailed numerical values for these metrics. The results demonstrate that the conditional PM RLHF approach outperforms the KL RLHF across various values of $\alpha$ and $\beta$. Specifically, this metric consistently yields lower values compared to the KL RLHF model, highlighting the efficacy of the conditional PM RLHF approach. Notably, when $\beta$ is set to 1.0, the conditional PM RLHF model with an $\alpha$ value of 0.3 exhibits better performance than the KL RLHF model, with a PM divergence of 1.57 compared to 2.23. Across all tested $\beta$ values, the optimal range for the $\alpha$ parameter appears to be 0.1 or 0.3 to achieve the minimal PM divergence.
Figure 6: The CDF of PM divergences across different RLHF Settings when fine-tuning the Llama-2-7B model. These figures illustrate the variability of PM divergences under four RLHF settings: Conditional PM RLHF with $\alpha$=0.1, $\alpha$=0.3, $\alpha$=0.5, and KL RLHF. Panels (a), (b), and (c) correspond to $\beta=0.1,0.5$, and $1$, respectively.
Distribution of PM Divergence.
Next, we examine the histogram and the cumulative distribution function (CDF) of instance preference matching divergence under various RLHF settings, as illustrated in Figure 5 and Figure 6. In Figure 5, panels (d), (e), and (f) compare the histogram of instance PM divergence for conditional PM RLHF (with the optimal $\alpha$) and KL RLHF. Across all $\beta$ values, the distribution of the instance PM divergence for models aligned using conditional PM RLHF is consistently closer to zero compared to those aligned with the KL RLHF method. This finding suggests that the conditional PM RLHF approach yields a higher concentration of instances with lower PM divergence values, indicating better alignment with the target preferences. In Figure 6, we compare the CDF of four different RLHF setups: conditional PM RLHF with $\alpha$ values of 0.1, 0.3, and 0.5, and a KL RLHF model for comparison. The distribution of the instance PM divergence for models aligned with conditional PM RLHF is closer to zero compared to those aligned with KL RLHF.
Table 2: Comparison between RLHF and conditional PM RLHF, when fine-tuning the Llama-2-7B model. The symbol $\downarrow$ means smaller is better. KL RLHF is equivalent to conditional PM RLHF when $\alpha$ is set to 1.
PM Divergence $\downarrow$ Length Perplexity $\downarrow$ Entropy KL
$\beta=0.1$ $\alpha=0.1$ 1.33 256 6.17 0.9 0.11
$\alpha=0.3$ 1.37 246 5.74 0.82 0.08
$\alpha=0.5$ 1.40 207 5.55 0.88 0.06
KL RLHF 1.53 131 5.48 0.92 0.05
$\beta=0.5$ $\alpha=0.1$ 1.58 256 7.17 0.78 0.15
$\alpha=0.3$ 1.46 253 5.57 0.81 0.07
$\alpha=0.5$ 1.57 221 5.13 0.95 0.03
KL RLHF 1.98 115 4.91 1.01 0.01
$\beta=1.0$ $\alpha=0.1$ 1.71 256 10.73 0.56 0.22
$\alpha=0.3$ 1.57 254 5.64 0.79 0.07
$\alpha=0.5$ 1.62 227 5.08 0.96 0.03
KL RLHF 2.23 108 4.82 1.05 0.01
Additional Measures.
Table 2 compares the performance of conditional PM RLHF and RLHF in terms of perplexity, length, entropy, and KL divergence when fine-tuning the Llama-2-7B model. We now examine the perplexity measure, which indicates a model’s predictive capabilities, with lower scores representing better performance. Across all $\beta$ values, as $\alpha$ increases, we observe a general decrease in perplexity, suggesting an improvement in the model’s language generation accuracy. At $\beta$ = 0.1, the perplexity decreases from 6.17 to 5.48 as we transition from conditional PM RLHF ($\alpha$=0.1) to RLHF, indicating better performance. This improving trend is consistent for $\beta$ = 0.5 and $\beta$ = 1.0. These trends suggest a trade-off between preference matching, as indicated by the PM divergence, and perplexity. However, maintaining $\alpha$ at or above 0.3 ensures that perplexity remains below 6, demonstrating good generation capabilities in LLMs. Consequently, the optimal balance between PM divergence and perplexity appears to be achieved with an $\alpha$ value in the range of 0.3 to 0.5. Within this range, LLMs attain an ideal equilibrium, effectively aligning with user preferences while preserving high-quality language generation.
@@ 6.2 Fine-Tuning LLMs
7 Discussion
In this paper, we have introduced PM RLHF to address algorithmic biases arising from aligning LLMs using a reward model. An integral component of this extension of RLHF is the PM regularizer, which takes the form of the negative logarithm of the probability distribution of the response. Essentially being the Shannon entropy of the LLM’s response, this regularization, together with the reward model, strikes a balance between reward maximization and response diversification. In particular, it allows the LLM to output responses that match the preference probabilities of the reward model under the PL model, thereby eliminating the algorithmic bias that standard RLHF suffers from. When applying PM RLHF to language domains, we resolve the challenge of the generated text becoming unnatural by introducing a conditional variant of PM RLHF. The conditional PM RLHF operates by excluding responses that have a very low probability with respect to a reference model. Our experimental results on pretrained Llama-2-7B and OPT-1.3B show that PM RLHF effectively reduces biases in aligning LLMs with the reward model’s preferences.
@@ 7 Discussion
To conclude the paper, we present several directions for future research. An interesting problem is to incorporate the PM regularization into using multiple reward models to model human preferences (Chakraborty et al., 2024; Zhong et al., 2024). This would help reduce algorithmic bias when human preferences exhibit heterogeneity. From a statistical perspective, an important direction is to examine whether the PM regularization in Theorem 3 under generalized PL models would offer practical benefits in preference alignment. Another important direction is to develop a DPO counterpart of PM RLHF. Although evidence is accumulating that DPO is less effective than RLHF for training highly performant LLMs, the low complexity and computational cost of implementing DPO, which owes to its bypassing of the reward learning step, render it popular for aligning small LLMs or for downstream tasks (Xiong et al., 2023). More broadly, an important question is to investigate how the length of the response would impact the reward as well as the preference (Park et al., 2024). It could possibly further reduce algorithmic biases using PM RLHF if the nuances of the length effect could be taken into account.
Acknowledgments
This work was supported in part by NIH grants, RF1AG063481 and U01CA274576, NSF DMS-2310679, a Meta Faculty Research Award, and Wharton AI for Business. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
Appendix A Additional Discussions of Algorithmic Bias
A.1 More Details on the PL Model
We start from providing the standard assumption of PL Model.
Assumption A.1 (The PL Model (Plackett, 1975; Luce, 2012)).
where
Here $y_{\tau(i)}\succ y_{\tau(1)},y_{\tau(2)},\dots,y_{\tau(i-1)},y_{\tau(i+1)},\ldots,y_{\tau(k)}$ represents the preference on $y_{\tau(i)}$ over the other responses. The PL model can be reduced to the BTL model in (2.1) by setting $\tau(k)=1$ and $\tau(k-1)=2$. In fact, a policy $\pi(y|x)$ that aligns with the PL model corresponds to aligning with the subsequent more refined probability:
which recovers Assumption A.1.
A.2 Bias from the Reward Model
In addition to the algorithmic bias of RLHF, we will also discuss the potential data bias and reward model bias induced by Step 2. In the data collection phase, a labeling team is responsible for ranking various sets of data and collect comparison information. However, it is essential to acknowledge that this team, like any other group, might have inherent biases. Their collective perspectives, backgrounds, and experiences might not perfectly represent the diverse range of human preferences. Given the collecting the data, the reward model might either underfit or overfit the comparison data. It can lead the agent to adopt behaviors that are not truly optimal or even potentially harmful.
Proposition A.2.
Proof: Given $x,y_{\textnormal{w}},y_{\textnormal{l}}$, denote
Here we omit the condition on $x,y_{\textnormal{w}},y_{\textnormal{l}}$, i.e., $p_{\textnormal{reward}^{\star}}(y_{\textnormal{w}}|x,y_{\textnormal{w}},y_{\textnormal{l}})$, to simplify the notation. Similarly, denote
Then, the objective function can be written as:
A.3 Bias from Beta
A.4 Other Discussions
Generating Examples for PM RLHF.
Human: What was the name of the movie that won the Academy Award for Best Picture in 1955?
Assistant: The movie that
Human: Why do birds migrate south for the winter?
Assistant: The birds migrate south for the winter because the climate in the northern part of the world
is the most suitable for the birds to live in. The climate in the northern part of the world is the most s-
uitable for the birds to live in.
Human: thank thank thank thank thank thank thank thank thank thank thank thank thank thank tha-
nk thank thank thank thank thank thank thank thank thank thank thank thank thank thank thank
Abbreviation Conditional Probability Formulation (i=$1,2$)
$p_{\text{rlhf}}(y_{i})$ $:=p_{\text{rlhf}}(y_{i}|y_{1},y_{2},x)$ $=\pi_{\text{rlhf}}(y_{i}|x)/(\pi_{\text{rlhf}}(y_{1}|x)+\pi_{\text{rlhf}}(y_{2}|x))$
$p_{\textnormal{ref}}(y_{i})$ $:=p_{\textnormal{ref}}(y_{i}|y_{1},y_{2},x)$ $=\pi_{\textnormal{ref}}(y_{i}|x)/(\pi_{\textnormal{ref}}(y_{1}|x)+\pi_{\textnormal{ref}}(y_{2}|x))$
$p_{\textnormal{reward}}(y_{i})$ $:=p_{\textnormal{reward}}(y_{i}|y_{1},y_{2},x)$ $=\exp(r(x,y_{i})/\beta)/(\exp(r(x,y_{1})/\beta)+\exp(r(x,y_{2})/\beta))$
Appendix B Proofs of Technical Results
B.1 Proof of Theorem 3
Recall that each $x$, assume that there are $k$ possible responses, namely, $y_{1},\ldots,y_{k}$. Since we can write
$\mathbb{E}_{y\sim\pi(\cdot|x)}\bigg{[}r(x,y)+R(\pi(y|x))\bigg{]}=\operatorname{\mathbb{E}}_{x}\operatorname{\mathbb{E}}_{y\sim\pi(\cdot|x)}\bigg{[}r(x,y)+R(\pi(y|x))\bigg{|}x\bigg{]},$
for a fixed $x$, without loss of generality, we assume the distribution $\pi(\cdot|x)$ is a multinomial distribution on $y_{1},\ldots,y_{k}$ with probabilities $\pi_{1},\ldots,\pi_{k}$, respectively. In addition, write $r_{i}=r(x,y_{i})$. Then it would be sufficient to find the the optimal $R$ to maximize
$\operatorname{\mathbb{E}}_{y\sim\pi(\cdot|x)}\bigg{[}r(x,y)+R\left(\pi(y|x)\right)\bigg{]}.$ (B.1)
By Hypothesis 3.2, the optimization problem above can be exactly minimized, which therefore is equivalent to maximizing
$\max_{\pi_{1},\ldots,\pi_{k}}\sum_{i=1}^{k}\pi_{i}(r_{i}+R(\pi_{i}))$
[…]
Assuming that $R$ is sufficiently differentiable, we next find the form of $R$ that leads to preference matching regularization. Now consider the Lagrangian
$L=\sum_{i=1}^{k}\pi_{i}(r_{i}+R(\pi_{i}))+\lambda\left(\sum_{i=1}^{k}\pi_{i}-1\right).$
The first-order conditions require that
$\frac{\partial L}{\partial\pi_{i}}=r_{i}+R(\pi_{i})+\pi_{i}R^{\prime}(\pi_{i})+\lambda=0,$
which must be satisfied by the unbiased policy $\pi^{\star}$, leading to
$R(\pi_{i}^{\star})+\pi_{i}^{\star}R^{\prime}(\pi_{i}^{\star})=-r_{i}-\lambda.$ (B.2)
Note that the preference matching probability $\pi^{\star}_{i}=\frac{g(r_{i})}{\sum_{j}g(r_{j})}$, which gives $r_{i}=g^{-1}(\pi_{i}^{\star})+c$, where $c$ does not depend on $i=1,\ldots,k$. Hence, from (B.2) we get
$R(\pi_{i}^{\star})+\pi_{i}^{\star}R^{\prime}(\pi_{i}^{\star})=-g^{-1}(\pi_{i}^{\star})+c^{\prime},$
where $c^{\prime}=-c-\lambda$ does not depend on $i$ either. Taking $\pi_{i}^{\star}$ as a variable as it in principle can take any value in $[0,1]$ and taking differentiation on both sides, we show that it must satisfy the following second-order ordinary differential equation:
$\pi R^{\prime\prime}(\pi)+2R^{\prime}(\pi)+\frac{1}{g^{\prime}(g^{-1}(\pi))}=0,$
which we call the preference matching differential equation in generalized PL model. Now we solve for the equation. Note that $pR^{\prime\prime}+2R^{\prime}=(pR)^{\prime\prime}$. Hence, we get
$R(\pi)=\frac{1}{\pi}\iint\pi R^{\prime\prime}+2R^{\prime}\mathrm{d}\pi=\frac{1}{\pi}\iint-\frac{1}{g^{\prime}(g^{-1}(\pi))}\mathrm{d}\pi=-g^{-1}(\pi)+\frac{G(g^{-1}(\pi))}{\pi}+C_{1}+\frac{C_{2}}{\pi},$
where $C_{1}$ and $C_{2}$ are constants.
In words, $\pi^{\star}(y|x)\equiv\mathrm{e}^{r(x,y)}/\sum_{y^{\prime}}\mathrm{e}^{r(x,y^{\prime})}$ is the global solution to (B.1) if $R$ takes the form above.
B.2 Proof of Theorem 4
@@
B.4 Proof of Proposition 5.1
The objective function in (2.3) is equal to

### v2

@@ B.2 Extreme Case: Preference Collapse in KL RLHF
The algorithmic bias and preference collapse stems from the use of a reference model in the RL objective (B.1) and, in particular, continues to occur even if the KL divergence is replaced by any $f$-divergence. The discussion of RLHF under $f$-divergence regularization is provided in the appendix.
To conclude this section, we summarize the key differences between PM RLHF and KL (or general $f$-divergence) RLHF. Both the reward model and the reference model implicitly encode preferences—the former aligned with human preferences, and the latter not. KL RLHF blends the preferences of these two models, which can lead to preference collapse. In contrast, PM RLHF focuses on extracting the preference encoded in the reward model while eliminating the influence of the reference model, thereby preserving human preferences.
Appendix C Other Related Work
Diversity in Human Preferences.
Training Algorithms.
Li et al. (2023b) demonstrated that PPO does not fully exploit the properties of RLHF in LLMs for preference alignment. Munos et al. (2023) introduced the Nash equilibrium of the preference model for aligning LLMs with human preferences, while Chen et al. (2024b) proposed a self-play-based approach to learn from supervised data as an alternative to supervised fine-tuning. Tang et al. (2024) discussed the performance gap between online and offline algorithms for alignment, and Ye et al. (2024) studied an online iterative RLHF algorithm with a general preference model. Outside of fine-tuning, model editing has emerged as a complementary strategy to modify LLM behavior across tasks (Jin et al., 2025).
A popular alternative to reward-model-based RLHF is direct preference optimization (DPO) (Rafailov et al., 2023), which directly fine-tunes LLMs on human preference data and has several notable variants (Liu et al., 2023; Azar et al., 2024; Chang et al., 2024; Gorbatovski et al., 2024; Rafailov et al., 2024; Yang et al., 2024). DPO is computationally more accessible since it bypasses the need for training a reward model. However, recent studies (Li et al., 2023a; Xu et al., 2024a; Tajwar et al., 2024) suggest that DPO is inferior to reward-model-based RLHF in aligning LLMs. Li et al. (2023a) and Xu et al. (2024a) empirically argued that representation misspecification is the reason why DPO is inferior to RL-based methods such as PPO. Furthermore, the on-policy nature of fine-tuning using a reward model can help improve LLM performance, possibly due to the distribution shift between the offline preference dataset during training and online responses during evaluation (Tajwar et al., 2024). The calibration challenges introduced by RLHF have also been studied by Xiao et al. (2025a). An alternative line of work introduces NLHF, which aims to align LLMs under general preference models (Munos et al., 2023). In this direction, Wang et al. (2025) examined the convergence behavior of NLHF and showed that the last iterate converges to a Nash equilibrium.
Appendix D Other Experiments
D.1 Perplexity of PM RLHF
When applying PM RLHF in practice, we encounter a numerical issue that is inherent to the nature of text data. A direct application of PM RLHF in our experiments fails to produce human-like text. This numerical issue corresponds to high values of perplexity, as illustrated by the experimental results in Figure 6. In particular, we evaluate the perplexity of the PM RLHF–aligned and KL RLHF–aligned models on 12,500 test samples. For the OPT-1.3B model (Zhang et al., 2022) fine-tuned by PM RLHF, the perplexity of the LLM is 2343.56, while KL RLHF yields a significantly lower perplexity of 7.24. Similarly, for the Llama-2-7B model (Touvron et al., 2023), the perplexity scores for LLMs trained by PM RLHF and KL RLHF are 7.15 and 5.48, respectively.
D.2 Fine-tuning OPT-1.3B Model
@@ D.1 Perplexity of PM RLHF
D.3 Other Discussion
@@
To conclude the paper, we present several directions for future research. An interesting problem is to incorporate the PM regularization into using multiple reward models to model human preferences (Chakraborty et al., 2024; Zhong et al., 2024). This would help reduce algorithmic bias when human preferences exhibit heterogeneity. Another important direction is to develop a DPO counterpart of PM RLHF. Although evidence is accumulating that DPO is less effective than RLHF for training highly performant LLMs, the low complexity and computational cost of implementing DPO, which is due to its ability to bypass the reward learning step, render it popular for aligning small LLMs or for downstream tasks (Xiong et al., 2023). More broadly, an important question is to investigate how the length of the response would impact the reward as well as the preference (Park et al., 2024). It could possibly further reduce algorithmic biases using PM RLHF if the nuances of the length effect could be taken into account.
Appendix E Proofs of Technical Results
E.1 Proof of Theorem 1
The discussion before the statement of Theorem 1 shows the necessity of (3.8). We finish the proof below by showing the sufficiency of (3.8). The objective function is:
$\mathbb{E}_{y\sim\pi(\cdot|x)}r(x,y)-\log(\pi(y|x)).$
$=$ $-D_{\text{KL}}(\pi(y|x)\|\pi^{\star}(y|x))+\mathbb{E}_{x}\log Z(x).$
It is also independent of $\pi(y|x)$. Therefore, we obtain
E.2 Proof of Theorem 4
@@
E.4 Proof of Proposition 5.1
The objective function in (A.2) is equal to